A vending machine costs Php 75,000 if paid in cash. On installment plan, a purchaser should pay Php 40,000 down payment and 8 quarterly installments, the first due at the BEGINNING of the first year. If money is worth 15% compunded monthly, determine the quarterly installment
Answer #1

Effective quarterly rate, r = (1 + 0.15/12)^4 - 1

Set the calculator to BGN mode

I/Y = 5.094533691

N = 8

PV = 75,000 - 40,000 = 35,000

FV = 0

CPT PMT

PMT = -5,172.463367

Quarterly installment = $5,172.463367

Can you please upvote? Thank You :-)